# The twisted coarse Baum–Connes conjecture and relative hyperbolic groups

### v1

@@ 2.2 The twisted coarse Baum–Connes conjecture
Assume that families of isometries $\{W_{1}(t)\}$ and $\{W_{2}(t)\}$ are covers of $f$. Since the unitary
$U(t):=\begin{pmatrix}1-W_{1}(t)W_{1}(t)^{*}&W_{1}(t)W_{2}(t)^{*}\\ W_{2}(t)W_{1}(t)^{*}&1-W_{2}(t)W_{2}(t)^{*}\\ \end{pmatrix}$
intertwines $\begin{pmatrix}W_{1}(t)TW_{1}(t)^{*}&0\\ 0&0\\ \end{pmatrix}$ and $\begin{pmatrix}0&0\\ 0&W_{2}(t)TW_{2}(t)^{*}\\ \end{pmatrix}$, it suffices to show that the function $\{U(t)\}$ is in the multiplier of $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))\otimes M_{2}$. It is easy to see that $\{W_{i}(t)W_{j}(t)^{*}\}$ is in the multiplier of $C^{*}_{L}(N)$ for $i,j=1,2$. For every partial translation $v$ on $Y$, $(T_{t})\in C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$ and $t\in[1,\infty)$, it suffices to show that the map $\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}:Y\to\mathcal{A}\otimes\mathcal{K}$ is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. We have that
$\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}(y)$ $=\chi_{C_{y}}\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)\chi_{C_{v}(y)}$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}W_{i}(t)W_{j}(t)^{*}\chi_{C_{v}(y)}\right)$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}(W^{\prime}_{i}(t)W^{\prime}_{j}(t)^{*}\otimes\text{id}_{\mathcal{A}})\chi_{C_{v(y)}}\right)$
for some covers $\{W_{i}^{\prime}(t)\}$ and $\{W_{j}^{\prime}(t)\}$ between Hilbert spaces $\ell^{2}(M_{0},\mathcal{H})$ and $\ell^{2}(N_{0},\mathcal{H})$. By the third condition in Definition 2.1, it is easy to see that $\left(T_{t}W_{i}(t)W_{j}(t)^{*}\right)^{v}$ is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. ∎
@@
Let $X$ and $Y$ be metric spaces with bounded geometry such that there exists a coarse equivalence $f:X\to Y$ and $g:Y\to X$. For any coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, if $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is the coarse $X$-algebra defined by the formula (4) along the coarse equivalence $f$, then we have that
@@
$V_{f}(t)$ $:\ell^{2}(P_{s}(X)_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(P_{s^{\prime}}(Y)_{0},\mathcal{H}_{\mathcal{A}})$
@@
$a$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i})(u)u^{*}$
$b$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i+1})(u)u^{*}$
@@
Combining this with Theorem 2.14, we obtain the following permanence result.
@@
Let $X$ and $Y$ be metric space with bounded geometry which are coarsely equivalent to each other. Then $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any coarse algebras if and only if $Y$ does.
@@
Moreover, if $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras then $Y$ does.
@@
Let $M$ be any metric space with bounded geometry and a net $X$ and a subset $Y\subset X$. Let $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ be any stable coarse $Y$-algebra and $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ are defined in the proof of Theorem 2.15. We define a $C^{*}$-algebra $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to be the closure of the *-algebra consisting all $(T_{t})_{t}$ in $\mathbb{C}_{L}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ such that for any partial translation $v$ there exists $r>0$ such that $(T_{t})^{v}\in\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ for all $t\in[1,\infty)$.
@@
induced by the natural inclusion $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\to C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$.
@@
We prove this by the induction on the dimension $m$ of $M$. When $m=0$, in this case we may assume $M=X$ and it suffices to show that the inclusion
@@
induces an isomorphism on $K$-theory, where for any $C^{*}$-algebra $\mathcal{B}$, $C_{ub}([1,\infty),\mathcal{B})$ is the $C^{*}$-algebra of all uniformly continuous bounded function from $1,\infty)$ to $\mathcal{B}$. But since $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K}))$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ are quasi-stable, the evaluation maps at 11
@@
induces isomorphisms on $K$-theory by [WY20, Lemma 12.4.3]. The assertion follows from the continuity of $K$-theory. Next, we assume induction hypothesis for the dimension $0,\cdots,m-1$. For each $m$-dimensional simplicial complex $\Delta$, we denote by $c(\Delta)$ the center of it. We define
@@
By taking a point from each simplex, we fix a net $X^{\prime}_{1}\subset N_{1}^{\prime}$, $X^{\prime}_{2}\subset N_{2}^{\prime}$ and $X_{1,2}^{\prime}\subset N^{\prime}_{1}\cap N^{\prime}_{2}$, respectively. Since $N_{1}^{\prime}$, $N_{2}^{\prime}$ and $N_{1}^{\prime}\cap N_{2}^{\prime}$ are strongly Lipschitz homotopic to simplicial complexes $N_{1}$, $N_{2}$ and $N_{1,2}$ whose dimensions are less than $m$, respectively through coarse equivalent maps. We denote the corresponding nets $X_{1}\subset N_{1}$, $X_{2}\subset N_{2}$ and $X_{1,2}\subset N_{1,2}$, respectively. By the same argument as Lemma 2.22, one can show that
$K_{*}(C^{*}_{L,\text{uniform}}(N_{i}^{\prime};\Gamma(X_{i}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L,\text{uniform}}(N_{i};\Gamma(X_{i},\mathcal{A}\otimes\mathcal{K})))$
@@
Note that $K$-theory for both $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ and $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ admits Mayer-Vietoris sequence by the same proof as [Yu97, Proposition 3.11]. Therefore the assertion follows from the five lemma and inductive hypothesis. ∎
@@
Take any unitary in the unitaization of $C^{*}_{L,\text{uniform}}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. We may assume that there exists $R\geq 0$ and a scalar $\lambda\in\mathbb{C}$ such that $u_{t}-\lambda$ is supported in $(P_{s}(N_{R}(Y)))\times(P_{s}(N_{R}(Y)))$ for every $t$. Therefore we can regard $u_{t}]$ as an element in $K_{*}(C^{*}_{L}(P_{s}(N_{R}(Y));\Gamma(Y,\mathcal{A}\otimes\mathcal{K})))$. But since we have an isomorphism
@@ 2.4 Family of subspaces
In this subsection, for a given metric space $X$ with bounded geometry and a family of subspaces $\{Y^{(j)}\}$ of $X$, we study the connection between the twisted coarse Baum–Connes conjecture of the family $\{Y^{(j)}\}$ and that of the space $X$.
We first realize the $K$-theory of twisted Roe algebras of $\bigsqcup_{j}Y^{(j)}$ as that of $X$ by changing the coefficient algebra $\mathcal{A}$. We denote by $\tilde{d}$ the distance on $\bigsqcup_{j}Y^{(j)}$ defined by
$\tilde{d}(x,y)=\left\{\begin{array}[]{cc}d(x,y)&x,y\in Y^{(j)}\text{ for some }j\\ \infty&\text{otherwise.}\end{array}\right$.
@@
$C^{*}(\bigsqcup_{j}X;\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}))\cong C^{*}(\bigsqcup_{j}Y^{(j)};\Gamma(\bigsqcup_{j}Y^{(j)},\mathcal{A}\otimes\mathcal{K})),$
@@
Let $\mathcal{A}$ be a $C^{*}$-algebra, $M$ a metric space with bounded geometry and a countable dense subset $M_{0}$ and a net $X\subset M$. For isometries $V:\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, and $W:\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, we denote the isometry induced on the geometric modules
$\tilde{V}:\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$ $\to\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$
@@
For any stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, the adjoint by $\tilde{V}$ and $\tilde{W}$ can be restricted to the twisted Roe algebra $C^{*}(M,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ by the fourth condition of Definition 2.1 and the induced map on $K$-theory is the identity map by the same arguements in the proof as [WY20, Lemma 5.1.12].
@@
Given a coarse $\bigsqcup_{j}X$-algebra $\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K})$, we define a new stable coarse $X$-algebra with respect to $\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K}$ as
@@
where $\psi:\mathcal{A}\otimes\mathcal{K}\otimes\mathcal{K}\cong\mathcal{A}\otimes\mathcal{K}$ is the fixed isomorphism. Then we have an isomorphism for Roe algebras and localization algebras on $K$-theory:
@@
$\alpha:$ $C^{*}\left(\bigsqcup_{j}X;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\to C^{*}\left(X;\Gamma(X,\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right)$
@@
$\beta(S)^{(j)}_{x,y}$ $:=\psi(S_{x,y}(j)).$
The compositions $\beta\circ\alpha$ and $\alpha\circ\beta$ are the adjoints by $\tilde{V}$ and $\tilde{W}$ in the previous remark, respectively, and both induce isomorphisms on $K$-theory. The same proof works for localization algebras. ∎

### v2

@@ 2.2 The twisted coarse Baum–Connes conjecture
Assume that families of isometries $\{W_{t}^{1}\}$ and $\{W_{t}^{2}\}$ are covers of $f$. Since the unitary
$U(t):=\begin{pmatrix}1-W_{t}^{1}(W_{t}^{1})^{*}&W_{t}^{1}(W_{t}^{2})^{*}\\ W_{t}^{2}(W_{t}^{1})^{*}&1-W_{t}^{2}(W_{t}^{2})^{*}\end{pmatrix}$
intertwines
$\begin{pmatrix}W_{t}^{1}T(W_{t}^{1})^{*}&0\\ 0&0\end{pmatrix}\quad\text{and}\quad\begin{pmatrix}0&0\\ 0&W_{t}^{2}(W_{t}^{2})^{*}\end{pmatrix}$,
it suffices to show that the function $\{U(t)\}$ is in the multiplier of $C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))\otimes M_{2}$. It is easy to see that $\{W_{t}^{i}(W_{t}^{j})^{*}\}$ is in the multiplier of $C^{*}_{L}(N)$ for $i,j=1,2$.
For every partial translation $v$ on $Y$, $(T_{t})\in C^{*}_{L}(N;\Gamma(Y,\mathcal{A}\otimes\mathcal{K}))$ and $t\in[1,\infty)$, it suffices to show that the map
$\left(T_{t}W_{t}^{i}(W_{t}^{j})^{*}\right)^{v}:Y\to\mathcal{A}\otimes\mathcal{K}$
is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. We have that
$\left(T_{t}W_{t}^{i}(W_{t}^{j})^{*}\right)^{v}(y)$ $=\chi_{C_{y}}\left(T_{t}W_{t}^{i}(W_{t}^{j})^{*}\right)\chi_{C_{v(y)}}$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}W_{t}^{i}(W_{t}^{j})^{*}\chi_{C_{v(y)}}\right)$
$=\sum_{z}\left(\chi_{C_{y}}T_{t}\chi_{C_{z}}\right)\left(\chi_{C_{z}}\bigl(\tilde{W}_{t}^{i}(\tilde{W}_{t}^{j})^{*}\otimes\mathrm{id}_{\mathcal{A}}\bigr)\chi_{C_{v(y)}}\right).$
for some covers $\{\tilde{W}_{t}^{i}\}$ and $\{\tilde{W}_{t}^{j}\}$ between Hilbert spaces $\ell^{2}(M_{0},\mathcal{H})$ and $\ell^{2}(N_{0},\mathcal{H})$. By the third condition in Definition 2.1, we obtain that
$\left(T_{t}W_{t}^{i}(W_{t}^{j})^{*}\right)^{v}$
is in $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$. This finishes the proof. ∎
@@
Let $X$ and $Y$ be metric spaces with bounded geometry. Assume that there exist coarse equivalences $f:X\to Y$ and $g:Y\to X$ such that the compositions $f\circ g$ and $g\circ f$ are close to the identity maps ${\rm id}_{Y}$ and ${\rm id}_{X}$, respectively. For any coarse $Y$-algebra $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$, if $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ is the coarse $X$-algebra induced the coarse equivalence $f$ via the formula (4), then we have that
@@
$V_{f}(t)$ $:\ell^{2}(P_{s}(X)_{0},\mathcal{H}_{\mathcal{A}})\to\ell^{2}(P_{s^{\prime}}(Y)_{0},\mathcal{H}_{\mathcal{A}}),$
@@
$a$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i})(u)u^{*},$
$b$ $=\bigoplus_{i=0}^{\infty}\text{ad}(W_{i+1})(u)u^{*},$
@@
Following [YU97, Proposition 3.7], these are in the unitaization of $C^{*}_{L}(P_{s^{\prime\prime}}(X))$. In addition, since
$a(t)-I=\text{diag}(\text{ad}(W_{0})(u)u^{*}-I,\cdots,\text{ad}(W_{N})(u)u^{*}-I,0,0\cdots),$
@@
Combining the above result with Theorem 2.14, we obtain the following permanence result.
@@
Let $X$ and $Y$ be metric spaces with bounded geometry. Assume that $X$ and $Y$ are coarsely equivalent. Then $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any coarse algebras if and only if $Y$ does.
@@
Moreover, if $X$ satisfies the twisted coarse Baum–Connes conjecture with respect to any stable coarse algebras, then so does $Y$.
@@
Let $M$ be any metric space with bounded geometry containing a net $X$, and $Y\subset X$ a subset of $X$. Let $\Gamma(Y,\mathcal{A}\otimes\mathcal{K})$ be any stable coarse $Y$-algebra and $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$ are defined in the proof of Theorem 2.15. We define a $C^{*}$-algebra $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ to be the closure of the *-algebra consisting all $(T_{t})_{t}$ in $\mathbb{C}_{L}[M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})]$ such that for any partial translation $v$ there exists $r>0$ such that $(T_{t})^{v}\in\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K})$ for all $t\in[1,\infty)$.
@@
induced by the natural inclusion
$C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K}))\to C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})).$
@@
We prove this by the induction on the dimension $m$ of $M$. In the case when $m=0$, we may assume $M=X$ and it suffices to show that the inclusion
@@
induces an isomorphism on $K$-theory, where for any $C^{*}$-algebra $\mathcal{B}$, $C_{ub}([1,\infty),\mathcal{B})$ is the $C^{*}$-algebra of all uniformly continuous, bounded function from $1,\infty)$ to $\mathcal{B}$. But since $\Gamma(N_{r}(Y),\mathcal{A}\otimes\mathcal{K}))$ and $\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ are quasi-stable, the evaluation-at-one maps
@@
induce isomorphisms on $K$-theory by [WY20, Lemma 12.4.3]. The assertion follows from the continuity of $K$-theory. Next, we assume induction hypothesis for the dimension $0,\cdots,m-1$. For each $m$-dimensional simplicial complex $\Delta$, we denote by $c(\Delta)$ the center of it. We define
@@
Taking a point from each simplex, we fix a net $X^{\prime}_{1}\subset N_{1}^{\prime}$, $X^{\prime}_{2}\subset N_{2}^{\prime}$ and $X_{1,2}^{\prime}\subset N^{\prime}_{1}\cap N^{\prime}_{2}$, respectively. Since $N_{1}^{\prime}$, $N_{2}^{\prime}$ and $N_{1}^{\prime}\cap N_{2}^{\prime}$ are strongly Lipschitz homotopic to simplicial complexes $N_{1}$, $N_{2}$ and $N_{1,2}$ whose dimensions are less than $m$, respectively through coarse equivalent maps. We denote the corresponding nets $X_{1}\subset N_{1}$, $X_{2}\subset N_{2}$ and $X_{1,2}\subset N_{1,2}$, respectively. By the same argument as Lemma 2.22, one can show that
$K_{*}(C^{*}_{L,\text{uniform}}(N_{i}^{\prime};\Gamma(X_{i}^{\prime},\mathcal{A}\otimes\mathcal{K})))$ $\cong K_{*}(C^{*}_{L,\text{uniform}}(N_{i};\Gamma(X_{i},\mathcal{A}\otimes\mathcal{K}))),$
@@
Note that $K$-theory for both $C^{*}_{L,\text{uniform}}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ and $C^{*}_{L}(M;\Gamma(X,\mathcal{A}\otimes\mathcal{K})))$ admits Mayer-Vietoris sequence by the same proof as [YU97, Proposition 3.11]. Therefore, the assertion follows from the five lemma and inductive hypothesis. ∎
@@
Take any unitary in the unitaization of $C^{*}_{L,\text{uniform}}(P_{s}(X);\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$. We may assume that there exists $R\geq 0$ and a scalar $\lambda\in\mathbb{C}$ such that $u_{t}-\lambda$ is supported in $(P_{s}(N_{R}(Y)))\times(P_{s}(N_{R}(Y)))$ for every $t$. Therefore we can regard $u_{t}]$ as an element in $K_{*}(C^{*}_{L}(P_{s}(N_{R}(Y));\Gamma(Y,\mathcal{A}\otimes\mathcal{K})))$. Since we have an isomorphism
@@ 2.4 Family of subspaces
In this subsection, for a given metric space $X$ with bounded geometry and a family of subspaces $\{Y^{(j)}\}_{j}$ of $X$, we study the connection between the twisted coarse Baum–Connes conjecture of the family $\{Y^{(j)}\}_{j}$ and that of the space $X$.
We first realize the $K$-theory of twisted Roe algebras of $\bigsqcup_{j}Y^{(j)}$ as that of $X$ by using a suitable coarse algebra. We denote by $\tilde{d}$ the distance on $\bigsqcup_{j}Y^{(j)}$ defined by
$\tilde{d}(x,y)=\left\{\begin{array}[]{cc}d(x,y),&\mbox{if~}x,y\in Y^{(j)}\text{ for some }j,\\ \infty,&\text{otherwise.}\end{array}\right$.
@@
$C^{*}\left(\bigsqcup_{j}X;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\cong C^{*}\left(\bigsqcup_{j}Y^{(j)};\Gamma\left(\bigsqcup_{j}Y^{(j)},\mathcal{A}\otimes\mathcal{K}\right)\right),$
@@
Let $\mathcal{A}$ be a $C^{*}$-algebra, $M$ a metric space with bounded geometry and a countable dense subset $M_{0}$ and a net $X\subset M$. For isometries $V:\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, and $W:\ell^{2}\otimes\mathcal{H}_{\mathcal{A}}\cong\mathcal{H}_{\mathcal{A}}\to\ell^{2}\otimes\mathcal{H}_{\mathcal{A}},~\xi\mapsto(\xi,0,0,\cdots)$, we denote the isometries induced on the geometric modules
$\tilde{V}:\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}})$ $\to\ell^{2}(M_{0},\mathcal{H}_{\mathcal{A}}),$
@@
For any stable coarse $X$-algebra $\Gamma(X,\mathcal{A}\otimes\mathcal{K})$, the adjoint by $\tilde{V}$ and $\tilde{W}$ can be restricted to the twisted Roe algebra $C^{*}(M,\Gamma(X,\mathcal{A}\otimes\mathcal{K}))$ by the fourth condition of Definition 2.1 and the induced map on $K$-theory is the identity map by the same argument in the proof as [WY20, Lemma 5.1.12].
@@
Let $M$ be a metric space with bounded geometry and fix a net $X\subset M$. Given a coarse $\bigsqcup_{j}X$-algebra $\Gamma(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K})$, we define a new stable coarse $X$-algebra with respect to $\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K}$ as
@@
where $\psi:\mathcal{A}\otimes\mathcal{K}\otimes\mathcal{K}\xrightarrow{\cong}\mathcal{A}\otimes\mathcal{K}$ is the fixed isomorphism. Then we have an isomorphism for Roe algebras and localization algebras on $K$-theory:
@@
$\alpha:$ $C^{*}\left(\bigsqcup_{j}X;\Gamma\left(\bigsqcup_{j}X,\mathcal{A}\otimes\mathcal{K}\right)\right)\to C^{*}\left(X;\Gamma(X,\ell^{\infty}(\mathbb{N},\mathcal{A}\otimes\mathcal{K})\otimes\mathcal{K})\right),$
@@
$\beta(S)^{(j)}_{x,y}$ $:=\psi(S_{x,y}(j)),$
where $e_{1,1}\in\mathcal{K}$ is any rank one projection. The compositions $\beta\circ\alpha$ and $\alpha\circ\beta$ are the adjoints by $\tilde{V}$ and $\tilde{W}$ in the previous remark, respectively, and both induce isomorphisms on $K$-theory. The same proof works for localization algebras. ∎